OpenStudy (anonymous):

Meta-math \[\int_{0}^{\infty} \text{J}_0(a\sqrt{1+x^2}) \ \text{d}x\]

OpenStudy (anonymous):

What is J_o

OpenStudy (anonymous):

@mukushla can u PLZ explain me the symbols?

OpenStudy (anonymous):

\(J_0\) is the bessel function of order 0

OpenStudy (experimentx):

looks like you are up to something!!

OpenStudy (experimentx):

\[ J_\alpha(x) = \sum_{m=0}^\infty \frac{(-1)^m}{m! \, \Gamma(m+\alpha+1)} {\left(\tfrac{1}{2}x\right)}^{2m+\alpha} \] \[ \int_0^\infty \sum_{m=0}^\infty \frac{(-1)^m}{m! \, \Gamma(m+1)}\left( {1 \over 2} a \sqrt{ 1 + x^2}\right)^{2m} \\ \]

OpenStudy (experimentx):